OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

use addition or elimination and add the two equations together you will get: 3x=15 then divide by 3 to get that x=5 plug in x=5 into the first equation to get: 5-y=13 and solve for y -y=8 so y=-8 check answer using the second equation: 2(5)+ (-8) = 2 10-8=2, 2=2 check so the answer is (5, -8)